OpenStudy (anonymous):

A rectangular field is two times as long as it is wide. If the perimeter of the field is 600, what are the dimensions of the field? Write and eqation using only 1 variable(w) to answer the given question. Let w be the width of the field.

OpenStudy (whpalmer4):

The answer to this is just like the second way I wrote the previous problem. Can you do it?

OpenStudy (mathstudent55):

Let L = length Let W = width two times as long as it is wide: L = 2W perimeter of the field is 600: P = L + L + W + W = 2L + 2W = 600, so 2L + 2W = 600, but L = 2W, so 2(2W) + 2W = 600 4W + 2W = 600 6W = 600

OpenStudy (anonymous):

Can you rewrite the way you did the other problem please?

OpenStudy (anonymous):

Oh thanks @mathstudent55

OpenStudy (whpalmer4):

Writing it terms of W: Perimeter = 2*length + 2*width, but length is 2*width P = 2*(2*W) + 2*W = 6W = 600 W = 100 L = 2*W = ?

OpenStudy (anonymous):

So L=200 and W=100?

OpenStudy (whpalmer4):

Bingo.

OpenStudy (whpalmer4):

Checking the result: 100 + 100 + 200 + 200 = 600

OpenStudy (anonymous):

Thank you so much. You're awesome :D